# On Reasoning Strength Planning in Large Reasoning Models

### v1

@@ 2.3 Activation Steering with Linear Direction
Recent advances in mechanism explainability reveal that there exist linear directions inside the activation space of language models that control specific semantical behaviors [33, 34, 35]. These directions are typically derived from activation differences between contrasting semantics, such as refusal versus compliance in responses [35]. Manipulating directions via activation addition or subtraction enables behavior modification of language models during inference. It has been evidenced that response style [25, 36], refusal behaviors [35, 37], and memory extraction capabilities [38] have been encoded in linear directions. Recent studies find evidence that by concatenating question prompts with their corresponding chain-of-thought (CoT) answers and computing activation differences between responses of varying reasoning token numbers, it is possible to identify linear directions that control the length of reasoning [19, 39, 40]. In this work, we take an important step further, by demonstrating that such directions have been pre-allocated by LRMs upon observing the question for reasoning strength control, even before generating the answer.
@@ 4.1 Pre-allocated Direction Vectors Exist for Reasoning Strength Planning
where the first and second terms denote the mean activations at layer $l$ computed over the activation sets $\mathcal{D}_{i}$ and $\mathcal{D}_{0}$, which correspond to math questions of difficulty level $i$ and 00, respectively. Here, these activations are also extracted at the start-of-reasoning token <think> position before generation. By varying the target difficulty $i$ from 1 to 5, we can get four such vectors at each layer, namely $\mathbf{r}^{(l)}_{5\leftarrow\text{1}}$, $\mathbf{r}^{(l)}_{4\leftarrow\text{1}}$, $\mathbf{r}^{(l)}_{3\leftarrow\text{1}}$, and $\mathbf{r}^{(l)}_{2\leftarrow\text{1}}$. These vectors capture the activation shift from a baseline level of difficulty 1 to increasingly harder questions. These vectors are pre-allocated since we extract them before generation. If the LRM plans the reasoning strength via a shared directional vector, then the four vectors are expected to show a high degree of similarity.
@@ 4.2 Pre-allocation Vectors Causally Affect the Reasoning Strengths
To further examine whether these direction vectors are used for reasoning strength planning, we test their causal effect on reasoning strengths via intervention of activation steering [25]. The key idea of activation steering is to inject direction vectors on the activation of language models, to test whether such direction vectors can causally affect the model behaviors [25, 35]. We conduct such activation steering experiments with the average vector $\mathbf{r}^{(l)}$ of our extracted four vectors (i.e., $\mathbf{r}^{(l)}=\frac{1}{4}\sum_{i=2}^{5}\mathbf{r}^{(l)}_{i\leftarrow\text{1}}$) as:
@@
[…]
@@ 4.4 Pre-allocation Vectors Control Reasoning Strengths by Modifying Logits of </think>
These pre-allocation vectors control the reasoning strength by modifying the logits of the end-of-reasoning token </think>. We visualize the distribution of logits for the </think> token under different steering strengths in Figure 8. As shown, applying a negative steering strength results in an overall increase in the logits of the </think> token, indicating a higher likelihood of its occurrence, which leads to fewer reasoning tokens. Conversely, positive steering strength decreases the logits of the </think> token, reducing the likelihood of generating this token, which leads to more reasoning token numbers. We also compare the impact on the logits with the eos token <|endoftext|> and randomly selected tokens. As shown in Figure 8(b), the impact on the end-of-reasoning token is significantly higher than random tokens and the eos token |endoftext|. These observations suggest that the pre-allocated vectors primarily modulate reasoning strength by adjusting the logits of the end-of-reasoning token. More similar results are in Appendix C.4.
@@ 7 Conclusion
This paper investigated whether and how large reasoning models (LRMs) plan reasoning strength (i.e., the number of reasoning tokens) before generating answers. Using linear probing, we showed that the reasoning token number can be predicted merely using the activations of the question, suggesting implicit reasoning strength planning capabilities before generation. We found the existence of pre-allocated direction vectors in LRMs, whose magnitude causally affects the number of reasoning tokens. We further revealed that these vectors may affect the logits of the end-of-thinking token </think> to achieve reasoning strength control. Finally, we discussed two potential applications of our findings: overthink detection before generation and efficient inference.
@@ D.2.1 Details
D.2.2 Case Study
@@
Question: What is the correct answer to this question: Kant’s humanity formulation of the categorical imperative makes it impermissible for us to, he argues, Choices: (A) sell one of our teeth. (B) engage in prostitution. (C) both A and B (D) Neither A nor B Only put the option letter in the box, e.g. boxedA. There is only one correct answer, .   $w/o$ Steering:<think> Okay, so I need to figure out what Kant’s humanity formulation says about selling one of our teeth or engaging in prostitution. Let me start by recalling what the humanity formulation is. From what I remember, Kant’s categorical imperative has several formulations, and one of them is the humanity formulation. The key idea here is that we should treat humanity, whether in oneself or in others, always as an end and never merely as a means to an end. So, the question is asking which actions Kant would consider impermissible under this principle. The options are selling one’s tooth, prostitution, both, or neither. First, let me break down each option. Selling a tooth: that’s essentially selling part of your body. Now, in the context of Kant’s ethics, using oneself as a means to an end might be problematic. $\cdots$ Wait, but what about the actual examples? I remember that Kant specifically mentions prostitution in his works, right? He considered it immoral because it treats the person as an object for sexual use, which is a means to an end. Similarly, selling body parts might be similar. But selling a tooth is a physical part, whereas prostitution is an action. However, in both cases, the person is using their body for monetary gain, possibly reducing themselves to a means. But maybe Kant would say that selling a tooth is permissible if it’s for a good reason, like saving a life, but not sure. Wait, the formulation is about treating humanity as an end. So any action that uses a person (including oneself) merely as a means would be wrong. So selling your tooth would involve using your body as a means to get money, but does that mean you’re treating yourself as a mere means? Maybe. Because you’re using your body as a commodity, rather than valuing it as an end in itself. Alternatively, maybe Kant would distinguish between selling a tooth (a part) and selling services (like prostitution). But I think both involve commodifying the body. So perhaps both are impermissible. Thus, the answer would be (C) both A and B. $\cdots$ Alternatively, maybe he only opposes prostitution, but I think both. Let me think again. Another angle: the humanity formulation requires that in any action, you never use a person merely as a means. So if you sell your tooth, you are using yourself as a means (to get money), but you’re also an end. However, if you are selling a tooth, maybe you are treating yourself as an end in some way. Wait, perhaps the distinction is whether you are using your body as a mere means. If you sell a tooth, you are using that part of your body as a tool, but does that mean you are treating your humanity as a means? The humanity is your rational nature. So maybe selling a tooth doesn’t directly violate that, as it’s a physical part. Alternatively, Kant might argue that the body is part of what makes a person, so using it as a commodity reduces oneself. So perhaps both are wrong. Since I’m a bit torn here, but I think the standard interpretation is that Kant would say both are impermissible. Hence, answer (C). </think>. (1610 tokens)   $\phantom{/o}w$ Steering:<think> "Okay, so I need to figure out what the answer is here. The question is about Kant’s humanity formulation of the categorical imperative and what he says is impermissible. The options are selling a tooth, prostitution, both, or neither. First, I remember that Kant’s categorical imperative has different formulations. $\cdots$ So both selling a tooth (if it’s your own) and prostitution would be impermissible because both involve treating humanity (either your own or someone else’s) as a means. So the answer should be C, both A and B? But wait, maybe selling a tooth isn’t considered the same as prostitution? Like, a tooth is part of your body but maybe it’s not using humanity as an end? Hmm, but Kant might say that even your own body shouldn’t be treated merely as a resource to exploit for gain. So selling a tooth could still be wrong because you’re using your own body as a means to an end (making money) rather than respecting your own humanity as an end in itself. So both A and B are impermissible. Therefore, the answer should be C. " </think> (418 tokens)
@@
Appendix E Broader Impacts

### v2

@@ 2.3 Activation Steering with Linear Direction
Activation steering is one research line among the representation learning [33, 34, 35, 36, 37, 38, 39]. Recent advances in mechanism explainability reveal that there exist linear directions inside the activation space of language models that control specific semantical behaviors [40, 41, 42]. These directions are typically derived from activation differences between contrasting semantics, such as refusal versus compliance in responses [42]. Manipulating directions via activation addition or subtraction enables behavior modification of language models during inference. It has been evidenced that response style [25, 43, 44], refusal behaviors [42, 45, 46], and memory extraction capabilities [47] have been encoded in linear directions. Recent studies find evidence that by concatenating question prompts with their corresponding chain-of-thought (CoT) answers and computing activation differences between responses of varying reasoning token numbers, it is possible to identify linear directions that control the length of reasoning [19, 48, 49]. In this work, we take an important step further, by demonstrating that such directions have been pre-allocated by LRMs upon observing the question for reasoning strength control, even before generating the answer.
@@ 4.1 Pre-allocated Direction Vectors Exist for Reasoning Strength Planning
where the first and second terms denote the mean activations at layer $l$ computed over the activation sets $\mathcal{D}_{i}$ and $\mathcal{D}_{0}$, which correspond to math questions of difficulty level $i$ and 0, respectively. Here, these activations are also extracted at the start-of-reasoning token <think> position before generation. By varying the target difficulty $i$ from 1 to 5, we can get four such vectors at each layer, namely $\mathbf{r}^{(l)}_{5\leftarrow\text{1}}$, $\mathbf{r}^{(l)}_{4\leftarrow\text{1}}$, $\mathbf{r}^{(l)}_{3\leftarrow\text{1}}$, and $\mathbf{r}^{(l)}_{2\leftarrow\text{1}}$. These vectors capture the activation shift from a baseline level of difficulty 1 to increasingly harder questions. These vectors are pre-allocated since we extract them before generation. If the LRM plans the reasoning strength via a shared directional vector, then the four vectors are expected to show a high degree of similarity.
@@ 4.2 Pre-allocation Vectors Causally Affect the Reasoning Strengths
[…]
@@
Table 1: Accuracy $\%$ comparison before and after steering across datasets. The performance is bold if improved after steering, and the absolute improvement is shown as superscripts.
@@ 4.4 Pre-allocation Vectors Control Reasoning Strengths by Modifying Logits of </think>
These pre-allocation vectors control the reasoning strength by modifying the logits of the end-of-reasoning token </think>. We visualize the distribution of logits for the </think> token under different steering strengths in Figure 8. As shown, applying a negative steering strength results in an overall increase in the logits of the </think> token, indicating a higher likelihood of its occurrence, which leads to fewer reasoning tokens. Conversely, positive steering strength decreases the logits of the </think> token, reducing the likelihood of generating this token, which leads to more reasoning token numbers. We also compare the impact on the logits with the eos token <|endoftext|> and randomly selected tokens. As shown in Figure 8(b), the impact on the end-of-reasoning token is significantly higher than random tokens and the eos token |endoftext|. These observations suggest that the pre-allocated vectors primarily modulate reasoning strength by adjusting the logits of the end-of-reasoning token. More similar results are in Appendix C.4. We also reveal that the steering mechanism also affects reasoning-related token logits within the reasoning process in Appendix C.5
@@ 7 Conclusion
This paper investigated whether and how large reasoning models (LRMs) plan reasoning strength (i.e., the number of reasoning tokens) before generating answers. Using linear probing, we showed that the reasoning token number can be predicted merely using the activations of the question, suggesting implicit reasoning strength planning capabilities before generation. We found the existence of pre-allocated direction vectors in LRMs, whose magnitude causally affects the number of reasoning tokens. We further revealed that these vectors may affect the logits of the end-of-thinking token </think> to achieve reasoning strength control. Finally, we discussed two potential applications of our findings: overthink detection before generation and efficient inference. Our paper studied the underlying mechanism of the reasoning strength planning in LRMs from the model activation perspective, which can help the community better understand LRMs  [^1] The broader impacts will be discussed in Appendix F.
Acknowledgments
This research/project is supported by the National Research Foundation, Singapore under its National Large Language Models Funding Initiative (AISG Award No: AISG-NMLP-2024-002). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore
@@ C.4 Pre-allocated Vectors Control Reasoning Strengths by Modifying Logits of </think>
C.5 Pre-allocated Vectors Control Reasoning-related Token Logits
We further test the changes in the logits of reasoning-related tokens. We find that with a positive steering strength, the logits of complex reasoning-related tokens also increase, beyond merely decreased logits of the end-of-reasoning token </think>.
We report the changes in the logits for these reasoning-related tokens as follows in Table 2. These tokens are usually regarded as reflection patterns in the R1-style response. Increasing the logits of these tokens increases the probabilities of reflection and increases of reasoning token numbers, and decreasing them has the opposite effect. As shown in Table 2, positive steering strengths increase the logits of these tokens while negative steering strengths decrease them. This indicates the reasoning strength control is not superficial.
Table 2: Impact of activation steering on logits of reasoning-related tokens.
Token $\lambda=-0.2$ $\lambda=-0.1$ $\lambda=0.1$ $\lambda=0.2$
Alright -0.2033 -0.0638 +0.0117 +0.0061
Hmm -1.1171e-4 -4.7589e-5 +2.5699e-5 +7.8185e-6
Oh -5.1966e-8 -2.9514e-8 +8.5295e-8 +2.2929e-7
Wait -7.3789e-9 -1.6405e-14 +2.8739e-8 +9.2974e-8
Additionally, to further test whether simply changing the logits of the end-of-reasoning token can achieve the same performance, we conduct experiments by multiplying the logits of the end-of-reasoning token with a factor $\gamma$ (i.e., , $\texttt{logit}_{new}=\gamma*\texttt{logit}_{old}$). We report the results on MATH500 in Table 3. We find that naively changing the logits will lead to unstable manners, either leading to overly increased answer token number (i.e., $\gamma$ = 2 on R1-Distill-Qwen-7B) or easily leading to decoding errors (i.e., $\gamma$ = 0.8).
These results further suggest that the changes in the logits of the end-of-reasoning token are the result of planning, rather than the cause of planning.
Table 3: Effect of changing the logits with $\gamma$.
Reasoning Token Number Answer Token Number Accuracy
R1-Distill-Qwen-1.5B
$\gamma$ = 2 4445.93 376.89 82.00
$\gamma$ = 1 4316.95 372.12 82.78
$\gamma$ = 0.8 16159.28 2.00 0.15
R1-Distill-Qwen-7B
$\gamma$ = 2 1013.48 5795.13 56.20
$\gamma$ = 1 3392.95 386.03 92.18
$\gamma$ = 0.8 11309.18 2.00 0.00
R1-Distill-Qwen-14B
$\gamma$ = 2 3267.30 387.06 94.35
$\gamma$ = 1 3148.91 391.72 93.37
$\gamma$ = 0.8 15913.95 2.00 0.15
R1-Distill-Qwen-32B
$\gamma$ = 2 3038.15 393.72 95.10
$\gamma$ = 1 3029.32 395.75 94.05
$\gamma$ = 0.8 6963.46 2.00 0.00
QwQ-32B
$\gamma$ = 2 3623.64 423.12 95.45
$\gamma$ = 1 3690.35 436.25 95.90
$\gamma$ = 0.8 16256.47 1.00 0.00
@@ D.2.1 Details
Appendix E Discussion
E.1 Generalization on More Domains
We discussed the generalization capabilities of our findings on other general domains beyond MATH, such as AlpacaEval and MMLU in Section 5, where our findings of reasoning strength prediction and strength control also hold on these two datasets. We also add one more experiment on the complex logic reasoning tasks such as GPQA diamond [9] and LiveCodeBench [64] to better demonstrate the generalization capabilities. As shown in Table 4 and Table 5, the number of reasoning tokens generally exhibits a similar pattern with the $\lambda$, indicating the generalization of this pre-allocation vector to other domains. Additionally, proper positive steering can also bring performance improvements. Since we found it hard to extract the generated code in a correct format for evaluation for code generation tasks when varying the response lengths, we just report the effect on the reasoning token numbers. There is only one exception on R1-Distill-Qwen-1.5B, where reducing the reasoning token number even brings the performance gain. We carefully analyze this model and find that when generating overlong responses, it tends to forget to follow the original instruction, and sometimes the results get cut off. Therefore, reducing the reasoning token number helps it avoid such bad issues like cut off, and increasing it will have the opposite effect.
Table 4: The effect of steering and performance on GPQA diamond.
$\lambda$
Model -0.2 -0.15 -0.1 -0.05 0 0.05 0.1 0.15 0.2
R1-Distill-Qwen-1.5B
Tokens 959.44 2930.29 5659.00 6677.72 7428.22 7690.48 7902.20 7805.64 7784.97
Accuracy 8.01 6.05 8.20 12.30 13.86 16.80 17.58 16.80 19.53
R1-Distill-Qwen-7B
Tokens 1022.10 2755.88 5250.92 6243.68 6998.01 7326.04 7875.52 8103.89 8380.32
Accuracy 5.66 12.30 22.27 28.13 33.92 36.13 40.23 39.26 40.33
R1-Distill-Qwen-14B
Tokens 886.26 2289.21 3910.04 6035.36 6365.84 6167.72 5927.13 5463.18 5137.19
Accuracy 15.42 17.97 23.24 44.34 50.78 51.95 51.56 51.76 50.20
R1-Distill-Qwen-32B
Tokens 266.07 296.68 1612.85 4696.14 5881.31 6391.68 6809.91 7128.37 7274.20
Accuracy 17.38 19.33 19.14 44.73 55.76 56.83 59.77 59.38 54.30
QwQ-32B
Tokens 6490.33 6826.64 7368.09 7472.53 8087.62 8811.63 8763.01 9161.22 9862.67
Accuracy 48.43 58.79 59.38 59.76 60.55 60.96 60.16 59.38 55.86
Table 5: The effect of steering and performance on LiveCodeBench (Code Execution).
$\lambda$
Model -0.2 -0.15 -0.1 -0.05 0 0.05 0.1 0.15 0.2
R1-Distill-Qwen-1.5B
Tokens 1971.90 2569.96 3112.51 3942.85 4682.46 5856.20 7120.98 8660.37 10482.23
Accuracy 26.30 23.02 18.32 13.20 8.72 5.48 2.30 1.41 0.84
R1-Distill-Qwen-7B
Tokens 281.09 358.46 678.20 1209.86 1625.67 2198.33 2942.55 3701.05 4688.05
Accuracy 61.85 66.75 72.65 79.33 79.23 79.30 74.63 80.17 80.85
R1-Distill-Qwen-14B
Tokens 346.57 461.33 929.89 1267.48 1428.35 1562.22 1745.25 1966.76 2197.66
Accuracy 68.42 74.01 82.05 91.54 91.03 91.28 89.35 89.35 89.67
R1-Distill-Qwen-32B
Tokens 230.60 345.78 501.81 827.44 1201.67 1682.29 2317.60 2916.74 3465.59
Accuracy 52.14 56.89 58.51 68.58 79.54 86.69 90.03 93.63 86.01
QwQ-32B
Tokens 1053.54 1151.99 1284.14 1439.79 1682.31 1895.04 2202.13 2521.79 2927.20
Accuracy 93.74 94.99 96.45 98.38 99.06 99.27 99.27 98.64 98.49
Table 6: The effect of steering and performance on LiveCodeBench (Code Generation).
$\lambda$
Model -0.2 -0.15 -0.1 -0.05 0 0.05 0.1 0.15 0.2
R1-Distill-Qwen-1.5B
Tokens 9176.89 9548.69 9915.46 9999.58 10422.80 10744.27 10963.86 11083.30 11070.14
R1-Distill-Qwen-7B
Tokens 1905.08 6409.04 8333.01 8747.09 8964.33 9194.85 9549.61 9635.65 10233.36
R1-Distill-Qwen-14B
Tokens 6732.97 6987.67 7578.29 7310.26 7299.31 7238.23 7008.38 6797.31 6494.35
R1-Distill-Qwen-32B
Tokens 3959.58 6786.27 6674.37 6735.45 6801.60 6969.60 7105.03 7702.86 8340.40
QwQ-32B
Tokens 5523.47 5563.86 5673.22 6040.58 6337.92 6921.26 7511.89 8164.05 8345.72
E.2 Generalization on More Model Backbones
We add experiments on another kind of LRM DeepSeek-R1-Distill-Llama-8B [2]. All of our findings also hold on this model, and we will add more results in our paper. We illustrate some datapoints of our main observations on DeepSeek-R1-Distill-Llama-8B [2]. We report the linear probing results in Table 7, Table 8, and Table 9. We also report the activation steering results in Table 10. Interestingly, DeepSeek-R1-Distill-Llama-8B is more sensitive to the steering strength $\lambda$.
Table 7: Prediction on middle layers.
Layer 25 26 27 28 29 30
R 0.8481 0.8467 0.8442 0.8464 0.8423 0.8430
Table 8: Cosine similarity between pre-allocated vectors on layer 31.
@@ E.2 Generalization on More Model Backbones
[…]
Table 9: Mean cosine similarities on middle layers.
Layer 25 26 27 28 29 30
R 0.9681 0.9661 0.8873 0.9620 0.9660 0.9124
Table 10: Activation Steering results across datasets.
$\lambda$
Dataset -0.02 -0.015 -0.01 -0.005 0 0.005 0.01 0.015 0.02
MATH500
Tokens 3042.42 3391.56 3312.67 3478.20 3457.01 3484.90 3530.19 3702.91 3654.19
Accuracy 75.18 80.00 81.78 83.55 82.80 83.38 83.98 84.88 84.48
AIME2024
Tokens 10366.34 10910.79 10458.87 10750.82 11049.53 10774.68 10888.95 11553.91 11622.77
Accuracy 26.67 32.08 34.78 38.75 38.96 38.75 37.08 42.50 41.67
OlympiadBench
Tokens 6926.34 7174.21 7254.21 7265.71 7234.64 7338.05 7522.05 7615.02 7345.39
Accuracy 43.69 48.24 49.94 50.13 49.76 49.59 50.41 51.20 49.30
E.2.1 Case Study
Question: Compute: $1-2+3-4+5-\cdots+99-100$.   $w/o$ Steering:<think>Okay, so I have this problem to compute the sum: 1 - 2 + 3 - 4 + 5 - 6 + … + 99 - 100. Hmm, that looks like an alternating series where the signs alternate between positive and negative. Let me try to figure this out step by step. First, let me write out the series to understand the pattern better: 1 - 2 + 3 - 4 + 5 - 6 + … + 99 - 100. I notice that this series alternates between adding an odd number and subtracting the next even number. So, each pair of terms is like (1 - 2), (3 - 4), (5 - 6), and so on. Maybe I can group them into these pairs and see what each pair adds up to. Let me compute a few pairs: 1 - 2 = -1, 3 - 4 = -1, 5 - 6 = -1, 7 - 8 = -1, … It seems like each pair of terms adds up to -1. That’s interesting. So, if I can figure out how many such pairs there are, I can multiply that number by -1 to get the total sum. Since the series starts at 1 and ends at 100, that’s a total of 100 numbers. If I’m pairing them two by two, the number of pairs should be 100 divided by 2, which is 50. So, there are 50 pairs.
$\cdots$
That is, for each term, if n is odd, we add it, and if n is even, we subtract it. So, the general term is $(-1)^{(n+1)}*n$. Maybe I can use a formula for such alternating series. I remember that the sum of an alternating series can sometimes be calculated using arithmetic series formulas or by grouping terms. Wait, another approach: Maybe I can separate the series into two separate sums: the sum of the odd terms and the sum of the even terms. Let’s see: The series is 1 + 3 + 5 + … + 99 minus (2 + 4 + 6 + … + 100). So, if I can compute both of these sums separately, I can subtract them to get the result. First, the sum of the odd numbers from 1 to 99. How many terms are there? Since we’re going up to 99, which is the 50th odd number (since 99 = 2*50 - 1). So, the number of terms is 50.
$\cdots$
[…]
</think>. (2253 tokens)
  $\phantom{/o}w$ Steering:<think>First, I observe that the sequence alternates between addition and subtraction of consecutive integers. To simplify the calculation, I can group the terms in pairs: (1 - 2), (3 - 4), (5 - 6), and so on, up to (99 - 100). Each pair results in -1. Since there are 100 numbers in total, there are 50 such pairs. Multiplying the number of pairs by each result gives 50 times -1, which equals -50.</think> (113 tokens)
@@ E.2.1 Case Study
Question: What is the correct answer to this question: Kant’s humanity formulation of the categorical imperative makes it impermissible for us to, he argues, Choices: (A) sell one of our teeth. (B) engage in prostitution. (C) both A and B (D) Neither A nor B Only put the option letter in the box, e.g. boxedA. There is only one correct answer, .   $w/o$ Steering:<think> Okay, so I need to figure out what Kant’s humanity formulation says about selling one of our teeth or engaging in prostitution. Let me start by recalling what the humanity formulation is. From what I remember, Kant’s categorical imperative has several formulations, and one of them is the humanity formulation. The key idea here is that we should treat humanity, whether in oneself or in others, always as an end and never merely as a means to an end. So, the question is asking which actions Kant would consider impermissible under this principle. The options are selling one’s tooth, prostitution, both, or neither. First, let me break down each option. Selling a tooth: that’s essentially selling part of your body. Now, in the context of Kant’s ethics, using oneself as a means to an end might be problematic.
$\cdots$
Wait, but what about the actual examples? I remember that Kant specifically mentions prostitution in his works, right? He considered it immoral because it treats the person as an object for sexual use, which is a means to an end. Similarly, selling body parts might be similar. But selling a tooth is a physical part, whereas prostitution is an action. However, in both cases, the person is using their body for monetary gain, possibly reducing themselves to a means. But maybe Kant would say that selling a tooth is permissible if it’s for a good reason, like saving a life, but not sure. Wait, the formulation is about treating humanity as an end. So any action that uses a person (including oneself) merely as a means would be wrong. So selling your tooth would involve using your body as a means to get money, but does that mean you’re treating yourself as a mere means? Maybe. Because you’re using your body as a commodity, rather than valuing it as an end in itself. Alternatively, maybe Kant would distinguish between selling a tooth (a part) and selling services (like prostitution). But I think both involve commodifying the body. So perhaps both are impermissible. Thus, the answer would be (C) both A and B.
$\cdots$
Alternatively, maybe he only opposes prostitution, but I think both. Let me think again. Another angle: the humanity formulation requires that in any action, you never use a person merely as a means. So if you sell your tooth, you are using yourself as a means (to get money), but you’re also an end. However, if you are selling a tooth, maybe you are treating yourself as an end in some way. Wait, perhaps the distinction is whether you are using your body as a mere means. If you sell a tooth, you are using that part of your body as a tool, but does that mean you are treating your humanity as a means? The humanity is your rational nature. So maybe selling a tooth doesn’t directly violate that, as it’s a physical part. Alternatively, Kant might argue that the body is part of what makes a person, so using it as a commodity reduces oneself. So perhaps both are wrong. Since I’m a bit torn here, but I think the standard interpretation is that Kant would say both are impermissible. Hence, answer (C).
</think>. (1610 tokens)
  $\phantom{/o}w$ Steering:<think> "Okay, so I need to figure out what the answer is here. The question is about Kant’s humanity formulation of the categorical imperative and what he says is impermissible. The options are selling a tooth, prostitution, both, or neither. First, I remember that Kant’s categorical imperative has different formulations.
$\cdots$
So both selling a tooth (if it’s your own) and prostitution would be impermissible because both involve treating humanity (either your own or someone else’s) as a means. So the answer should be C, both A and B? But wait, maybe selling a tooth isn’t considered the same as prostitution? Like, a tooth is part of your body but maybe it’s not using humanity as an end? Hmm, but Kant might say that even your own body shouldn’t be treated merely as a resource to exploit for gain. So selling a tooth could still be wrong because you’re using your own body as a means to an end (making money) rather than respecting your own humanity as an end in itself. So both A and B are impermissible. Therefore, the answer should be C. "
</think> (418 tokens)
@@
Appendix F Broader Impacts